Natasha inherits $50,000 and decides to give $20,000 to charity. With the remaining money she invests, into two stocks paying 16% and 25%. If she makes $6060 on the two investments, find how much money is invested in each stock. Please include variables and representation.
Answer by mananth(16946)

Investment I 16.00% per annum ---x
Investment II 25.00% per annum ---y

x + y= 30000 ------------------------1
16.00% x + 25.00% y = $6,060.00
Multiply by 100
16 x + 25 y = $606,000.00 --------2

Multiply (1) by -16
we get

-16 x -16 y = -480000.00

Add this to (2)

0 x 9 y = $126,000.00

divide by 9

y = $14,000.00 investment at 25.00%
Balance $16,000.00 investment at 16.00%
CHECK
$16,000.00 --------- 16.00% ------- $2,560.00
$14,000.00 ------- 25.00% ------- $3,500.00
Total ------- $6,060.00
m.ananth@hotmail.ca